# RESEARCH ARTICLE

### v3

@@ 1.  Introduction
Automatic Music Transcription (AMT), which converts audio signals into symbolic musical notation, represents a fundamental challenge in Music Information Retrieval (MIR) (Benetos et al.,, 2019). While current systems achieve remarkable accuracy in transcribing polyphonic music for specific instrument timbres (Hawthorne et al.,, 2017; Riley et al.,, 2024; Tamer et al., 2023a, ), we investigate a more practical scenario: processing audio mixtures containing $K$ instrument classes and outputting $K$ tracks of musical notes, each corresponding to a distinct timbre. We term this task “timbre-separated transcription”, which differs from conventional source separation by operating at the symbolic level to extract note representations rather than reconstructing audio waveforms. In this paper, we study the musical sounds produced by instruments, and the term “timbre” refers specifically to instrument categories (e.g., violin, viola, flute), distinguished from individual physical sources.
Current solutions often use unified classification-based models that treat timbre identification as a categorization task (Wu et al.,, 2020; Gardner et al.,, 2022), which limits their flexibility. They demand extensive training data, fix the maximum number of separable sources, and fail to generalize to unseen timbres, essentially functioning as a “timbre dictionary”. Moreover, these models are typically large and computationally demanding, making them inaccessible to general users despite advances in transcription accuracy.
To address these issues, we propose a lightweight two-branch architecture that decouples pitch/onset estimation from timbre representation learning and employs deep clustering (Hershey et al.,, 2016) to achieve timbre-separated transcription. Our key advance over prior deep clustering transcription (Tanaka et al.,, 2020) is note-level clustering (vs. frame-level), tailored for symbolic output. The first branch performs timbre-agnostic transcription, predicting frame-level activations with a compact, fully convolutional network. The second branch learns direction-aware timbre embeddings, which are clustered at the note level to enable dynamic instrument separation. Furthermore, we discuss the impact of dataset and possible improvements. All code is open-sourced, with the model already deployed in a web-based assistive transcription tool11endnote: 1Link to the web tool:
https://madderscientist.github.io/noteDigger/
@@ 2.1  AMT Fundamentals
Stage 1 is typically implemented with classical signal-processing transforms. The Short-Time Fourier Transform (STFT) yields linear-spaced frequency bins, whereas the Mel-scale Transform (Mel) and Constant-Q Transform (CQT) provide logarithmic spacing that better matches musical pitch perception. However, these interpretable transforms may not always be optimal, leading to explorations of learnable encodings (Luo and Mesgarani,, 2018; Ravanelli and Bengio,, 2018; Zeghidour et al.,, 2021) and hybrid methods (Su and Yang,, 2015; Rouard et al.,, 2023). The CQT is readily enriched into a Harmonic CQT (HCQT) (Bittner et al.,, 2017) by vertically shifting the spectrogram according to the bin offsets of successive harmonics and concatenating the shifted copies along an additional axis (Balhar and Hajiˇc,, 2019). This prior-driven expansion allows small convolutional kernels to focus efficiently on musically relevant frequency components.
Stage 2 produces a 2-D posteriogram that encodes the probability of pitch activation at each time-frequency bin. Convolutional layers leverage the spectro-temporal grid to capture local structures (Wu et al.,, 2019; Bittner et al.,, 2022), while temporal-sequence models extend the receptive field along time, modeling long-range dependencies (Hawthorne et al.,, 2017). To translate frame-wise posteriors into discrete notes, in stage 3, most systems rely on onset detection (i.e., the initial frame of each note) and then link successive active frames within an onset-defined segment to form complete note events (Hawthorne et al.,, 2017; Wu et al.,, 2020, 2024; Bittner et al.,, 2022).
@@ 2.2  Timbre-Separated Transcription
Multi-instrument timbre-separated transcription goes beyond traditional AMT by extracting instrument-specific notes from polyphonic mixtures. Tamer et al., 2023b demonstrated that simply cascading source separation followed by transcription leads to suboptimal results due to error propagation. Therefore, it’s necessary to combine transcription and separation in a mutually reinforcing manner.
Innovative methods have emerged for this challenge. Gardner et al., (2022) introduced a sequence-to-sequence AMT model that outputs multiple instrument tracks with explicit instrument assignments, achieving high accuracy but relying heavily on diverse multi-timbre data. Multi-task learning frameworks like Timbre-Trap (Cwitkowitz et al.,, 2024) and Cerberus (Manilow et al.,, 2020) show that joint optimization can lead to synergistic effects. Wu et al., (2020) proposed a self-attention-based instance segmentation approach that performs joint note detection and instrument classification, achieving timbre-separated transcription for a closed set of trained instruments.
For handling unseen timbres, researchers have explored different strategies. The zero-shot learning method proposed by Lin et al., (2021) employs a query-based mechanism: under contrastive supervision, it learns to encode a clean reference example into a timbre embedding that generalizes to unseen instruments. This embedding is then used to sequentially modulate the U-Net layers, enabling zero-shot separation and transcription of audio sources with novel timbres not encountered during training. The work by Tanaka et al., (2020) pioneered the application of deep clustering networks to timbre-separated transcription. Their approach develops a generalizable timbre encoder and enables separation with a non-fixed number of classes through clustering, showing how these tasks can complement each other when properly combined.
@@ 2.3  Deep Clustering Methodology
The pioneering work of Hershey et al., (2016) first applied deep clustering to separate unknown speakers. This was later extended by Luo et al., (2018), who introduced learnable attractors as speaker-specific reference points in an embedding space that cluster time-frequency bins, enabling end-to-end separation of a variable number of sources. The adaptation to music transcription was first achieved by Tanaka et al., (2020), but their method uses transcription mainly to aid audio separation, lacks dedicated note-creation post-processing, and omits key implementation details. More fundamentally, these studies share two critical limitations: clustering at frame level produces fragmented results, and the hard-assignment strategy cannot handle overlap situations. In contrast, our clustering post-process is tailored for music transcription, aggregating frame embeddings into coherent note events with minimal computational overhead.
@@ 3.2  Overall Architecture
The timbre-agnostic transcription branch builds upon BasicPitch (Bittner et al.,, 2022), a lightweight fully convolutional network that demonstrates strong performance using only local contextual information. We identify several limitations and introduce targeted optimizations, yielding a more compact architecture with comparable accuracy.
The timbre encoding branch adopts a similar fully convolutional design. To incorporate global context, we employ InstanceNorm and GroupNorm for adaptive normalization across the entire input. While convolutional architectures are more deployment-friendly on low-resource devices than RNN-based approaches commonly used in deep clustering (Hershey et al.,, 2016; Tanaka et al.,, 2020), they may lack long-range modeling capacity; alternative designs are discussed in the experiments.
@@ 3.3  EnergyNorm for Spectral Normalization
Conventional two‑dimensional normalization schemes often involve subtractive operations and compute statistics directly over the time–frequency plane. These practices alter the relative amplitude relationships among harmonic components within the same analysis frame, thereby undermining the acoustic basis of timbre perception.
@@ 3.5  Focal Loss for Class Imbalance
We directly utilized Focal Loss (Lin et al.,, 2017) in Eqn. (7) with $\gamma=1$, setting the positive class weights for notes and onsets at 0.2 and 0.06, respectively.
@@ 3.6  InfoNCE for Contrastive Cluster Formation
Hershey et al., (2016) and Tanaka et al., (2020) utilize the following deep clustering objective:
@@
This loss function benefits from low spatial complexity and high computational efficiency. However, it’s at least three orders of magnitude larger than the AMT loss (frame&onset transcription with Focal Loss) and converges slowly, which makes joint optimization more difficult when a backbone is shared. Furthermore, this objective enforces orthogonality between encodings of different classes. We argue that strict orthogonality is an overly stringent constraint; for clustering objectives, anti-parallel alignment yields larger inter-class distances than orthogonality. Consequently, we adopt the contrastive InfoNCE loss (Oord et al.,, 2018) to facilitate cluster formation:
@@
3.7  Post-Process for Timbre-Separated Transcription
@@
Tanaka et al., (2020) proposed a frame-level transcription method that predicts a binary mask by assigning an additional cluster to “silence bins” and performing K-means clustering over the entire time–frequency space. But this approach suffers from several critical limitations:
@@
  • Most clustering algorithms require specifying the number of clusters, and the hard-assignment strategy cannot handle overlap situations, where instruments play the same pitch simultaneously.
@@
For the last issue, we attempted an iterative matching-filter approach, which, however, tends to merge categories and can only be performed at the frame level, so will not be elaborated further.
@@ 4.1.1  Datasets
Table 1 summarizes the datasets used in our experiments. We train primarily on MusicNet (Thickstun et al.,, 2017), but replace its DTW-aligned annotations, which are known to be inaccurate, with the refined labels from MusicNetEM (Maman and Bermano,, 2022). All audio is resampled to 22,050 Hz and split into non-overlapping 900-frame, i.e., 10.4-second clips (frame size: 256 samples, matching the CQT hop length). Tracks sharing the same instrument are mixed to form a single timbral class per piece.
For evaluation, we use three real-world polyphonic datasets with solo tracks: BACH10 (Duan et al.,, 2010) (fixed quartet of violin, bassoon, clarinet, saxophone), URMP (Li et al.,, 2019), and PHENICX (Miron et al.,, 2016), the latter featuring complex orchestral mixtures with an average of 9.5 instrument classes per piece.
Motivated by Slakh2100 (Manilow et al.,, 2019), a large-scale fully synthesized dataset, we question the necessity of human-composed, real-recorded data. We therefore generate synthetic audio via a simple algorithm that randomly places notes across our 84-note pitch range and adds harmonically overlapping chord tones with probability. MIDI sequences are rendered with FluidSynth, applying randomization in dynamics, pitch tuning, and articulation to improve realism. Each sample uses a single timbre; an example appears in Figure 4.
@@
Table 2: Comparison of timbre-agnostic transcription. Unless stated otherwise, models are trained on our synthetic dataset. “noLog”: EnergyNorm without log; “BN”: BatchNorm replacing EnergyNorm (BasicPitch-style); “Conv39”: BasicPitch’s 39-tap conv replacing our dilated conv; “BPloss”: BasicPitch’s loss; “MN”: trained on MusicNet; lCQT”: learnable CQT; $\text{CQT}_{\text{MN}}: l$earnable CQT trained on MusicNet; “BP”: re-implementation of BasicPitch (Bittner et al.,, 2022); $\text{BP}^{\text{fl}}$: BP with focal loss; $\text{BP}^{\text{fl}}_{\text{MN}}$: $\text{BP}^{\text{fl}}$ trained on MusicNet; “OF”: re-implementation of Onsets&Frames (Hawthorne et al.,, 2017) with its original BCE loss; $\text{OF}^{\text{fl}}$: OF with focal loss; $\text{OF}^{\text{fl}}_{\text{MN}}$: $\text{OF}^{\text{fl}}$ trained on MusicNet. In headers, “tutti” = full-mix polyphonic pieces; “stems” = single-instrument samples.
@@
Table 3: Comparison of timbre-separated transcription. Unless stated otherwise, models are trained on MusicNet with InfoNCE loss (10). “D16”: $D=16$; “MSE”: using $L_{\text{affinity}}$(3.6); “Syn”: trained on our synthetic dataset; “Rescale”: forcibly scaling amplitude using Frame prediction before InstanceNorm in timbre encoding branch; “Share”: sharing the first residual block between two branches; “Tanaka”: baseline model (Tanaka et al.,, 2020) using InfoNCE. All experiments use identical pre-trained AMT branch parameters except “Share”.
@@ 4.1.2  Evaluation Metrics
We evaluate transcription performance at frame and note levels using mir_eval (Raffel et al.,, 2014). Since BasicPitch’s note creation requires thresholds for frames and onsets, we employ a coarse-to-fine search strategy to identify optimal thresholds that maximize model performance, assuming the metric is concave in the threshold. The search converges to $10^{-5}$ precision: first optimizing the frame threshold by maximizing frame-wise $F_{1}$ ($F_{F}$), then fixing it while tuning the onset threshold to maximize note-level $F_{1}$ ($F_{N}$). A note is correct if its pitch matches ground truth and its onset is within $\pm 50$ ms of the reference.
@@ 4.2.1  Parameter Setting
We use the parameter-efficient downsampling CQT architecture of Schörkhuber and Klapuri, (2010) with a hop size of 256 samples ($\sim$11 ms at 22,050 Hz). The timbre embedding dimension is set to $D=12$. Training and validation data follow a 10:1 ratio. We optimize with AdamW with initial $lr=3\times 10^{-4}$. Training lasts 60 epochs with a batch size of 18, and the model is selected based on the minimum validation loss. For timbre-separated transcription on our synthetic dataset, we generate mixtures of 2 or 3 timbres per sample and add mild white noise. The final model comprises: CQT (19,944 parameters), AMT branch (18,978), and timbre encoder (25,983).

### v4

@@ 1.  Introduction
Automatic Music Transcription (AMT), which converts audio signals into symbolic musical notation, represents a fundamental challenge in Music Information Retrieval (MIR) (Benetos et al., 2019). While current systems achieve remarkable accuracy in transcribing polyphonic music for specific instrument timbres (Hawthorne et al., 2017; Riley et al., 2024; Tamer et al., 2023a), we investigate a more practical scenario: processing audio mixtures containing $K$ instrument classes and outputting $K$ tracks of musical notes, each corresponding to a distinct timbre. We term this task “timbre-separated transcription”, which differs from conventional source separation by operating at the symbolic level to extract note representations rather than reconstructing audio waveforms. In this paper, we study the musical sounds produced by instruments, and the term “timbre” refers specifically to instrument categories (e.g., violin, viola, flute), distinguished from individual physical sources.
Current solutions often use unified classification-based models that treat timbre identification as a categorization task (Wu et al., 2020; Gardner et al., 2022), which limits their flexibility. They demand extensive training data, fix the maximum number of separable sources, and fail to generalize to unseen timbres, essentially functioning as a “timbre dictionary”. Moreover, these models are typically large and computationally demanding, making them inaccessible to general users despite advances in transcription accuracy.
To address these issues, we propose a lightweight two-branch architecture that decouples pitch/onset estimation from timbre representation learning and employs deep clustering (Hershey et al., 2016) to achieve timbre-separated transcription. Our key advance over prior deep clustering transcription (Tanaka et al., 2020) is note-level clustering (vs. frame-level), tailored for symbolic output. The first branch performs timbre-agnostic transcription, predicting frame-level activations with a compact, fully convolutional network. The second branch learns direction-aware timbre embeddings, which are clustered at the note level to enable dynamic instrument separation. Furthermore, we discuss the impact of dataset and possible improvements. All code is open-sourced, with the model already deployed in a web-based assistive transcription tool11endnote: 1Link to the web tool: https://madderscientist.github.io/noteDigger/.
@@ 2.1  AMT Fundamentals
Stage 1 is typically implemented with classical signal-processing transforms. The Short-Time Fourier Transform (STFT) yields linear-spaced frequency bins, whereas the Mel-scale Transform (Mel) and Constant-Q Transform (CQT) provide logarithmic spacing that better matches musical pitch perception. However, these interpretable transforms may not always be optimal, leading to explorations of learnable encodings (Luo and Mesgarani, 2018; Ravanelli and Bengio, 2018; Zeghidour et al., 2021) and hybrid methods (Su and Yang, 2015; Rouard et al., 2023). The CQT is readily enriched into a Harmonic CQT (HCQT) (Bittner et al., 2017) by vertically shifting the spectrogram according to the bin offsets of successive harmonics and concatenating the shifted copies along an additional axis (Balhar and Hajiˇc, 2019). This prior-driven expansion allows small convolutional kernels to focus efficiently on musically relevant frequency components.
Stage 2 produces a 2-D posteriogram that encodes the probability of pitch activation at each time-frequency bin. Convolutional layers leverage the spectro-temporal grid to capture local structures (Wu et al., 2019; Bittner et al., 2022), while temporal-sequence models extend the receptive field along time, modeling long-range dependencies (Hawthorne et al., 2017). To translate frame-wise posteriors into discrete notes, in stage 3, most systems rely on onset detection (i.e., the initial frame of each note) and then link successive active frames within an onset-defined segment to form complete note events (Hawthorne et al., 2017; Wu et al., 2020, 2024; Bittner et al., 2022).
@@ 2.2  Timbre-Separated Transcription
Multi-instrument timbre-separated transcription goes beyond traditional AMT by extracting instrument-specific notes from polyphonic mixtures. Tamer et al. (2023b) demonstrated that simply cascading source separation followed by transcription leads to suboptimal results due to error propagation. Therefore, it is necessary to combine transcription and separation in a mutually reinforcing manner.
Innovative methods have emerged for this challenge. Gardner et al. (2022) introduced a sequence-to-sequence AMT model that outputs multiple instrument tracks with explicit instrument assignments, achieving high accuracy but relying heavily on diverse multi-timbre data. Multi-task learning frameworks like Timbre-Trap (Cwitkowitz et al., 2024) and Cerberus (Manilow et al., 2020) show that joint optimization can lead to synergistic effects. Wu et al. (2020) proposed a self-attention-based instance segmentation approach that performs joint note detection and instrument classification, achieving timbre-separated transcription for a closed set of trained instruments.
For handling unseen timbres, researchers have explored different strategies. The zero-shot learning method proposed by Lin et al. (2021) employs a query-based mechanism: under contrastive supervision, it learns to encode a clean reference example into a timbre embedding that generalizes to unseen instruments. This embedding is then used to sequentially modulate the U-Net layers, enabling zero-shot separation and transcription of audio sources with novel timbres not encountered during training. The work by Tanaka et al. (2020) pioneered the application of deep clustering networks to timbre-separated transcription. Their approach develops a generalizable timbre encoder and enables separation with a non-fixed number of classes through clustering, showing how these tasks can complement each other when properly combined.
@@ 2.3  Deep Clustering Methodology
The pioneering work of Hershey et al. (2016) first applied deep clustering to separate unknown speakers. This was later extended by Luo et al. (2018), who introduced learnable attractors as speaker-specific reference points in an embedding space that cluster time-frequency bins, enabling end-to-end separation of a variable number of sources. The adaptation to music transcription was first achieved by Tanaka et al. (2020), but their method uses transcription mainly to aid audio separation, lacks dedicated note-creation post-processing, and omits key implementation details. More fundamentally, these studies share two critical limitations: clustering at frame level produces fragmented results, and the hard-assignment strategy cannot handle overlapping situations. In contrast, our clustering post-process is tailored for music transcription, aggregating frame embeddings into coherent note events with minimal computational overhead.
@@ 3.2  Overall Architecture
The timbre-agnostic transcription branch builds upon BasicPitch (Bittner et al., 2022), a lightweight fully convolutional network that demonstrates strong performance using only local contextual information. We identify several limitations and introduce targeted optimizations, yielding a more compact architecture with comparable accuracy.
The timbre encoding branch adopts a similar fully convolutional design. To incorporate global context, we employ InstanceNorm and GroupNorm for adaptive normalization across the entire input. While convolutional architectures are more deployment-friendly on low-resource devices than RNN-based approaches commonly used in deep clustering (Hershey et al., 2016; Tanaka et al., 2020), they may lack long-range modeling capacity; alternative designs are discussed in the experiments.
@@ 3.3  EnergyNorm for Spectral Normalization
Conventional two-dimensional normalization schemes often involve subtractive operations and compute statistics directly over the time–frequency plane. These practices alter the relative amplitude relationships among harmonic components within the same analysis frame, thereby undermining the acoustic basis of timbre perception.
@@ 3.5  Focal Loss for Class Imbalance
We directly utilized Focal Loss (Lin et al., 2017) in Eqn. (7) with $\gamma=1$, setting the positive class weights for notes and onsets at 0.2 and 0.06, respectively.
@@ 3.6  InfoNCE for Contrastive Cluster Formation
Hershey et al. (2016) and Tanaka et al. (2020) utilize the following deep clustering objective:
@@
This loss function benefits from low spatial complexity and high computational efficiency. However, it is at least three orders of magnitude larger than the AMT loss (frame&onset transcription with Focal Loss) and converges slowly, which makes joint optimization more difficult when a backbone is shared. Furthermore, this objective enforces orthogonality between encodings of different classes. We argue that strict orthogonality is an overly stringent constraint; for clustering objectives, anti-parallel alignment yields larger inter-class distances than orthogonality. Consequently, we adopt the contrastive InfoNCE loss (Oord et al., 2018) to facilitate cluster formation:
@@
3.7  Post-Processing for Timbre-Separated Transcription
@@
Tanaka et al. (2020) proposed a frame-level transcription method that predicts a binary mask by assigning an additional cluster to ”silence bins” and performing K-means clustering over the entire time–frequency space. But this approach suffers from several critical limitations:
@@
  • Most clustering algorithms require specifying the number of clusters, and the hard-assignment strategy cannot handle overlapping situations, where instruments play the same pitch simultaneously.
@@
The last issue remains unresolved, as our attempted iterative matching-filter approach suffers from category merging and is limited to frame-level processing; thus, we do not discuss it further.
@@ 4.1.1  Datasets
Table 1 summarizes the datasets used in our experiments. We train primarily on MusicNet (Thickstun et al., 2017), but replace its DTW-aligned annotations, which are known to be inaccurate, with the refined labels from MusicNetEM (Maman and Bermano, 2022). All audio is resampled to 22,050 Hz and split into non-overlapping 900-frame, i.e., 10.4-second clips (frame size: 256 samples, matching the CQT hop length). Tracks sharing the same instrument are mixed to form a single timbral class per piece.
For evaluation, we use three real-world polyphonic datasets with solo tracks: BACH10 (Duan et al., 2010) (fixed quartet of violin, bassoon, clarinet, saxophone), URMP (Li et al., 2019), and PHENICX (Miron et al., 2016), the latter featuring complex orchestral mixtures with an average of 9.5 instrument classes per piece.
Motivated by Slakh2100 (Manilow et al., 2019), a large-scale fully synthesized dataset, we question the necessity of human-composed, real-recorded data. We therefore generate synthetic audio via a simple algorithm that randomly places notes across our 84-note pitch range and adds harmonically overlapping chord tones with probability. MIDI sequences are rendered with FluidSynth, applying randomization in dynamics, pitch tuning, and articulation to improve realism. Each sample uses a single timbre; an example appears in Figure 4.
@@
Table 2: Comparison of timbre-agnostic transcription. Unless stated otherwise, models are trained on our synthetic dataset. “noLog”: EnergyNorm without log; “BN”: BatchNorm replacing EnergyNorm (BasicPitch-style); “Conv39”: BasicPitch’s 39-tap conv replacing our dilated conv; “BPloss”: BasicPitch’s loss; “MN”: trained on MusicNet; lCQT”: learnable CQT; $\text{CQT}_{\text{MN}}: l$earnable CQT trained on MusicNet; “BP”: re-implementation of BasicPitch (Bittner et al., 2022); $\text{BP}^{\text{fl}}$: BP with focal loss; $\text{BP}^{\text{fl}}_{\text{MN}}$: $\text{BP}^{\text{fl}}$ trained on MusicNet; “OF”: re-implementation of Onsets&Frames (Hawthorne et al., 2017) with its original BCE loss; $\text{OF}^{\text{fl}}$: OF with focal loss; $\text{OF}^{\text{fl}}_{\text{MN}}$: $\text{OF}^{\text{fl}}$ trained on MusicNet. In headers, “tutti” = full-mix polyphonic pieces; “stems” = single-instrument samples.
@@
Table 3: Comparison of timbre-separated transcription. Unless stated otherwise, models are trained on MusicNet with InfoNCE loss (10). “D16”: $D=16$; “MSE”: using $L_{\text{affinity}}$(3.6); “Syn”: trained on our synthetic dataset; “Rescale”: forcibly scaling amplitude using Frame prediction before InstanceNorm in timbre encoding branch; “Share”: sharing the first residual block between two branches; “Tanaka”: baseline model (Tanaka et al., 2020) using InfoNCE. All experiments use identical pre-trained AMT branch parameters except “Share”.
@@ 4.1.2  Evaluation Metrics
We evaluate transcription performance at frame and note levels using mir_eval (Raffel et al., 2014). Since BasicPitch’s note creation requires thresholds for frames and onsets, we employ a coarse-to-fine search strategy to identify optimal thresholds that maximize model performance, assuming the metric is concave in the threshold. The search converges to $10^{-5}$ precision: first optimizing the frame threshold by maximizing frame-wise $F_{1}$ ($F_{F}$), then fixing it while tuning the onset threshold to maximize note-level $F_{1}$ ($F_{N}$). A note is correct if its pitch matches ground truth and its onset is within $\pm 50$ ms of the reference.
@@ 4.2.1  Parameter Setting
We use the parameter-efficient downsampling CQT architecture of Schörkhuber and Klapuri (2010) with a hop size of 256 samples ($\sim$11 ms at 22,050 Hz). The timbre embedding dimension is set to $D=12$. Training and validation data follow a 10:1 ratio. We optimize with AdamW with initial $lr=3\times 10^{-4}$. Training lasts 60 epochs with a batch size of 18, and the model is selected based on the minimum validation loss. For timbre-separated transcription on our synthetic dataset, we generate mixtures of 2 or 3 timbres per sample and add mild white noise. The final model comprises: CQT (19,944 parameters), AMT branch (18,978), and timbre encoder (25,983).